# Collet–Eckmann maps in the unicritical family

### v1

@@ Abstract
In this paper we study perturbations of complex unicritical polynomials satisfying the Collet–Eckmann condition. We show that Collet–Eckmann parameters are Lebesgue density points of the complement of the Mandelbrot set (i.e. the connectedness locus).
@@ 1 Introduction
We consider the unicritical family
@@
where $d\geq 2$ is an integer and $c\in{\mathbb{C}}$.
A non-hyperbolic unicritical map $f_{c}$ satisfies the Collet–Eckmann condition (CE) if there exist $C>0$ and $\gamma>0$ such that
@@
Recall that a map $f$ in this family is hyperbolic if $\omega(0)\cap{\mathcal{J}}(f)=\emptyset$, where ${\mathcal{J}}(f)$ is the Julia set of $f$. For $d\geq 2$, the set $\mathcal{M}_{d}$ is defined as the set of parameters $c$ such that the Julia set ${\mathcal{J}}(f_{c})$ of $f_{c}$ is connected (or equivalently, the critical orbit $\{f^{n}_{c}(0)\}$ is bounded). When $d=2$, the set $\mathcal{M}_{2}={\mathcal{M}}$ is the well known Mandelbrot set (but we also call ${\mathcal{M}}_{d}$ a Mandelbrot set).
Our paper is about perturbations of CE-maps in the unicritical family. It has a quite long history. In [24], J. Rivera-Letelier proved that Misiurewicz maps (i.e. critically non-recurrent maps) are Lebesgue density points of ${\mathbb{C}}\setminus{\mathcal{M}}_{d}$. It was generalised by the first author in [1] to rational maps. However, even if Misiurewicz maps behave quite nicely under perturbations, they are rare in the parameter space ([2]).
In [6], A. Avila, M. Lyubich and W. Shen proved that the Lebesgue density of ${\mathcal{M}}_{d}$ at at most finitely renormalizable maps in the unicritical family, is strictly less than one (Theorem 1.4 and Remark 6.1 in [6]). Since Collet–Eckmann maps are at most finitely renormalizable, the result applies to them and consequently CE-maps cannot be density points of ${\mathcal{M}}_{d}$.
In a series of quite recent papers [12, 15, 16], J. Graczyk and G. Świa̧tek made careful estimates of the geometry of the Julia set around typical Collet–Eckmann maps with respect to harmonic measure on $\partial{\mathcal{M}}_{d}$. These maps are typically critically slowly recurrent, i.e. not Misiurewicz maps, but the speed of recurrence is slow (see [14]). Similar to the Misiurewicz case (which is easier), the goal is to get some kind of similarity between phase and parameter space. One cannot expect a complete (conformal) similarity between them since the Mandelbrot set contains an interior and the Julia set for a CE-map is a dendrite. However, at small scales this similarity becomes asymptotically conformal. The result is obtained by introducing a powerful tool called the similarity map, which transfers the geometry around the critical value $c$ between the Julia set and the Mandelbrot set ${\mathcal{M}}_{d}$ in an intriguing way. In particular it follows that typical Collet–Eckmann maps, with respect to harmonic measure, are density points of ${\mathbb{C}}\setminus{\mathcal{M}}_{d}$. The slow recurrence condition is also imposed on the Collet–Eckmann maps in [4], where we proved a corresponding result for rational maps. We also mention an unpublished work by M. Benedicks and J. Graczyk, where they prove corresponding results for one-dimensional families (with some regularity) around typical CE-maps with respect to harmonic measure in $\partial{\mathcal{M}}_{d}$. For more general CE-maps, the dynamical constants can be drastically different. This yields quite wild dynamics, and crucial parts of the traditional methods break down.
In [13] it is shown that at most finitely renormalizable maps are weak density points of ${\mathbb{C}}\setminus{\mathcal{M}}_{d}$ (again, this applies to CE-maps). The combinatorics for such maps seem to allow for the existence of deep central returns in the puzzle of the induced map at infinitely many small scales. Seemingly, this implies that this weak density cannot be replaced by strong density. We claim that this is not the case. Our main result is the following.
@@
One should compare the above theorem with a corresponding result in the real setting. M. Tsujii proved that (slowly recurrent) CE-maps with all periodic cycles repelling are Lebesgue density points of CE-maps [25] in families of real polynomials, including the unicritical family, restricted to ${\mathcal{M}}_{d}\cap{\mathbb{R}}$; see also [10].
It will be quite clear from the proof that the same methods used in this paper can be applied to get the same conclusion as in Theorem 1.1 for rational CE-maps with precisely one critical point in the Julia set.
Recall that a rational map satisfies the Collet–Eckmann condition if the map has no parabolic cycles and (1.1) is satisfied for each critical point in the Julia set (usually one requires the map to be non-hyperbolic). Let $\mathcal{R}_{d}$ be the space of rational maps of degree $d\geq 2$.
Theorem 1.3.
Any rational Collet–Eckmann map of degree $d$ at least two with only simple critical points and with precisely one critical point in the Julia set is a Lebesgue density point of hyperbolic maps in $\mathcal{R}_{d}$.
This also partially improves our previous result in [4] where we imposed the additional assumption of slow recurrence of the critical points in the Julia set.
Remark 1.4.
In case that rational maps have critical points of higher multiplicities, one can also get a corresponding result but in a subspace of $\mathcal{R}_{d}$. This subspace was defined by G. Levin [19].
Ideas of the proof.
We trace the critical orbits for all parameters close to a Collet–Eckmann map and use the parameter exclusion technique developed by M. Benedicks and L. Carleson in the seminal works [7, 8]. However, since we start with arbitrary Collet–Eckmann maps, the dynamical constants, e.g. the recurrence rate, can be quite wild in general (deep central returns) and traditional methods break down. This difficulty will be overcome by introducing the idea of promotion. Roughly speaking, promotion means that almost all parameters will reach some kind of slow recurrence condition.
Acknowledgement.
Partially this work was written when all the authors were participating in the program “Two Dimensional Maps” at Mittag-Leffler Institute during the spring 2023. We thank the organisers for providing a nice working place. We are also grateful to Michael Benedicks for stimulating discussions on the topic of this paper. The second author gratefully acknowledges support from the Knut and Alice Wallenberg Foundation. The third author was gratefully supported by Vergstiftelsen and the Qilu Young Scholar Program of Shandong University.
@@ 2 Preliminaries
Throughout the paper we fix $d\geq 2$. Let $f_{c_{0}}$ be a Collet–Eckmann unicritical polynomial of degree $d$. The only critical point is 00 with the corresponding critical value being $c_{0}$. Let ${\mathcal{Q}}={\mathcal{Q}}(c_{0},\varepsilon)$ be a square of side length $\varepsilon$ in the parameter plane centred at the parameter $c_{0}$ with $\varepsilon>0$ to be sufficiently small (cf. Lemma 2.9 and Lemma 2.8). For a parameter $c\in{\mathcal{Q}}$ the corresponding map is denoted, as usual, by $f_{c}$. We study the evolution of the critical orbit for nearby parameters. Let
@@ 2.1 Recurrence, bound periods and partitions
After the bound period ends, the two critical orbits will then separate and stay away from critical points until the next return happens. A return might be deep in the sense that $\xi_{n}(c)$ is quite close to 00. To achieve our goal, we will need to disregard parameters with very deep returns. For later purpose we will define $\alpha_{n}(c)$ for a parameter $c$ as follows:
@@
[…]
@@
For a partition element, we can also give a notion of returns. Let $A$ be a partition element at time $k-1$. Then $\xi_{k}(A)$ is called a return to $U$ if $\xi_{k}(A)\cap U\neq\emptyset$ and a return to $U^{\prime}$ if $\xi_{k}(A)\cap U^{\prime}\neq\emptyset$. It also may happen that $\xi_{k}(A)\cap U^{\prime}\neq\emptyset$ but $\xi_{k}(A)\cap U=\emptyset$. In this case we say that $\xi_{k}(A)$ is a pseudo-return. Now bound periods can also be defined similarly if a return occurs for a partition element.
@@ 2.2 Free periods
Let $f$ be rational satisfying the exponential shrinking property. Then there exist a neighbourhood $U^{\prime}$ of $\operatorname{Crit}(f)\cap{\mathcal{J}}(f)$, $C>0$ and $\gamma>0$ such that the following holds: Let $z\in{\mathcal{J}}(f)$ and if $f^{j}(z)\not\in U^{\prime}$ for $10\leq j\leq n-$ 1, then
@@
The following lemma follows from the above and is proved in [4, Lemma 3.10]. To state the result, let $\mathcal{N}_{\varepsilon_{0}}$ be a neighbourhood of the Julia set ${\mathcal{J}}(f_{c_{0}})$, where $\varepsilon_{0}$ is chosen in a way such that $U^{\prime}\subset\mathcal{N}_{\varepsilon_{0}}$. That is,
@@ 2.3 Transversality
A direct consequence of this is the following lemma proved in [3, Proposition 4.1].
@@ 3.1 Distortion, length and expansion during bound periods
Suppose that $\operatorname{dist}(0,\xi_{n}(A))\sim e^{-r}$. Then for all $z\in\xi_{n}(A)$ we have $|f_{c}(z)-c|\sim e^{-dr}$. We divide the sum into two parts $[0,j-1]=[0,J]\cup[J+1,j-1]$, where $J=\lceil r/(\alpha+\overline{\gamma})\rceil$ and where $\alpha$ is the upper bound of $\alpha_{j}=\alpha_{j}(c)$ for $c\in A$, satisfying $\alpha<\overline{\gamma}-\underline{\gamma}$. Then the first sum becomes
@@ 3.2 Main Distortion Lemma
Let $\varepsilon^{\prime}>0$ and $\tilde{\gamma}>0$. Then there exists $N$ large enough such that the following holds: If $A\subset{\mathcal{E}}_{\nu}(\tilde{\gamma})$ is a partition element and $\nu\geq N$ is a return time or does not belong to the bound period, and $\nu^{\prime}$ is the next free return, then we have
@@
For consistency we include the proof of this important lemma, although its proof is almost the same, verbatim, as the proof of Lemma 3.14 in [4] (alternatively see [3, Lemma 7.3]).
@@
where we in the sum from $q_{t}$ to $n-1$ used Lemma 2.8 (the first inequality, now with respect to $U^{\prime}$) and that $|\xi_{j}(b)|>\delta^{\prime}>\delta$ during this time.
@@ 2. $(ii)$
    $\mathcal{Q}$ is a partition element at time $N-1$, and $\xi_{N}(Q)$ either is an essential return or reaches the large scale $S$.
@@ 4. $(iv)$
    There exists a constant $K=K(\delta,\varepsilon_{1})>1$ such that for all $c\in\mathcal{Q}$ and $10\leq k\leq N-$ 1
@@
By the definition of a partition element we can use Lemma 3.8 repeatedly starting from the time $N_{L}$ up to $N-1$ to get the distortion claimed in (iii).
@@ 4 Promotion
By Lemma 3.9, all critical orbits for parameters in ${\mathcal{Q}}$ will follow the starting critical orbit up to time $N$ with bounded distortion and exponential growth of derivatives. If ${\mathcal{Q}}$ has grown to the large scale under $\xi_{N}$, then $\xi_{N}({\mathcal{Q}})$ covers mostly the Fatou set of $f_{0}$, and by a compactness argument, we can conclude that most parameters in ${\mathcal{Q}}$ will land in the Fatou set of the corresponding map under $\xi_{n}$; see next section for more details. Hence a very large portion of parameters are hyperbolic in ${\mathcal{Q}}$. Therefore, we may assume, without loss of generality, that $\xi_{N}({\mathcal{Q}})$ has not reached the large scale yet and thus is an essential return.
@@
which can be made arbitrarily small if $N$ is large enough. We make the deletion so that the smallest collection of whole partition elements are deleted (this means that we may delete a little more than necessary). The remaining parameters are partitioned into partition elements as above. Each of these partition elements, containing $c$, is denoted by $A_{N}(c)$. Sometimes we just write $A_{N}$. However, note that $A_{N-1}={\mathcal{Q}}$ (there is no partition before time $N$).
After this time $N$ we lose this kind of dependence on the starting orbit. So we need to control recurrence rates so that we do not lose a lot of derivative growth.
Let $N=\nu_{0}<\nu_{1}<\cdots$ be the free returns after time $N$. Suppose that $A$ is a partition element at time $\nu_{j}$. We will introduce two sequences $\{\widetilde{\alpha}_{\nu_{j}}(A)\}$ and $\{\underline{\gamma}_{\nu_{j}}(A)\}$ inductively which capture two key features for later analysis. Firstly, parameters $c\in A$ for which $\xi_{\nu_{j}}(c)$ are at distance $e^{-\widetilde{\alpha}_{\nu_{j}}(A)\nu_{j}}$ away from the critical point will be kept (and thus has controlled approach rate to the critical point), and secondly, we do not lose a lot of the derivative growth between two free returns.
We start with the time $\nu_{0}$. By Lemma 3.9 (iv), $\xi_{k}(c)$ stays close to $\xi_{k}(c_{0})$ and thus $\alpha_{k}(c)$ is comparable to $\alpha_{k}(c_{0})$ for all $k\leq N$ and all $c\in A_{N}$.
Definition of $\widetilde{\alpha}_{\nu_{0}}(A)$.
@@
$\widetilde{\alpha}_{\nu_{0}}(A_{N})=\sup_{c\in A_{N}}\alpha_{\nu_{0}}(c).$
@@
Let $|\xi_{\nu_{0}}(c)|\sim e^{-r}$, for some $c\in A_{N}$, where $A_{N}\subset{\mathcal{Q}}$ is a partition element satisfying (4.1). Then $|Df_{c}^{p_{0}+1}(\xi_{\nu_{0}}(c))|\geq e^{(1-\kappa)r}$.
@@
We can make $C_{1}$ as close to $1$ as we want so that $c\in A_{N}$ implies $c\in{\mathcal{E}}_{n}(\underline{\gamma})$. The lemma now follows from Corollary 3.5. ∎
Note that by the definition of $\kappa$, Lemma 4.1 says that $\operatorname{diam}\left(\xi_{\nu_{0}+p_{0}}(A_{N})\right)\geq e^{-\kappa r}$, where $\operatorname{diam}\left(\xi_{\nu_{0}}(A_{N})\right)\sim e^{-r}$. Now we fix some $\kappa^{\prime}$ such that $\kappa<\kappa^{\prime}<1$. In the following, we shall define $\{\widetilde{\alpha}_{\nu_{j}}(A)\}$ and $\{\underline{\gamma}_{\nu_{j}}(A)\}$, where $A$ is a partition element at the free return time $\nu_{j}$.
@@
$\gamma_{n}(A_{n})=\inf\limits_{c\in A_{n}}\gamma_{n}(c).$
@@
$\underline{\gamma}_{\nu_{j+1}}(A)=\left(\underline{\gamma}_{\nu_{j}}(A)+\gamma_{H}\frac{\ell_{j}}{\nu_{j}}\right)\frac{\nu_{j}}{\nu_{j+1}-1},$ (4.6)
@@
Note that $\underline{\gamma}_{\nu_{j}}$ assigns a number for each partition element. So, if $A_{j}$ and $A_{j+1}$ are respectively partition elements at time $\nu_{j}$ and $\nu_{j+1}$ and moreover $A_{j+1}\subset A_{j}$, we have $\underline{\gamma}_{\nu_{j}}(A_{j+1})=\underline{\gamma}_{\nu_{j}}(A_{j})$.
@@
$\geq\left(\underline{\gamma}_{\nu_{j}-1}(A)\left(1-\frac{1}{\nu_{j}}\right)+(d-1)\alpha_{\nu_{j}}(c)\left(1-\frac{1}{\nu_{j}}\right)\right)\frac{\nu_{j}}{\nu_{j+1}-1}$ (4.11)
@@
$\geq\left(\underline{\gamma}_{\nu_{j}-1}(A)\left(1-\frac{1}{\nu_{j}}\right)+\gamma_{H}\frac{\ell_{j}}{\nu_{j}}-\frac{C}{\nu_{j}}\right)\frac{\nu_{j}}{\nu_{j+1}-1}$ (4.13)
@@
$\frac{\underline{\gamma}_{\nu_{j}-1}+2C}{\nu_{j}}\leq\left((d-1)\left(1-\frac{1}{\nu_{j}}-\frac{\kappa^{\prime}}{\tilde{\kappa}}\right)+(1-\kappa)\right)\alpha_{\nu_{j}}$
for sufficiently large $\nu_{j}\geq N$, since $\alpha_{\nu_{j}}(c)\nu_{j}=r_{\nu_{j}}\geq\Delta$ and $\Delta$ is large. Then, using that $p_{j}\leq\nu_{j}$ by Lemma 3.2, we get $\gamma_{\nu_{j+1}}(c)\geq\underline{\gamma}_{\nu_{j+1}}(A)$. If we define $\underline{\gamma}_{k}(A)=\underline{\gamma}_{\nu_{j+1}}(A)$ for $\nu_{j}<k\leq\nu_{j+1}$, then it follows that

### v2

@@ Abstract
In this paper we show that Collet–Eckmann parameters in the unicritical family $f_{c}(z)=z^{d}+c$, where $d$ is at least $2$, are Lebesgue density points of the complement of the Mandelbrot set.
Mathematics subject classifications: 37F10, 37F15, 37F46.
@@ 1 Introduction
Consider the unicritical family
@@
where $d\geq 2$ is an integer and $c\in{\mathbb{C}}$ is a parameter. This family has been a central object of study in complex dynamics for a long time. One of the main objectives is to understand it from a topological and metric point of view. On the topological side, the famous Fatou conjecture, stating that the set of hyperbolic parameters is open and dense, was settled for real parameters $c$, for $d=2$, by J. Graczyk and G. Świa̧tek [16], and independently M. Lyubich [27]. It was settled in full generality, that is, for real polynomials, by O. Kozlovski, W. Shen and S. van Strien [22]. On the metric side, seminal works by M. Lyubich [28] and A. Avila and G. Moreira [8] together show that almost all real quadratic polynomials are either hyperbolic (nice robust maps with an attracting cycle) or expanding in a strong sense (i.e. satisfy the Collet–Eckmann condition, see below).
The Collet–Eckmann condition arose early as a natural condition for the presence of chaotic behaviour and the existence of absolutely continuous invariant measures (acim) [12, 14, 13]. The pioneering work of M. Jakobson [21] showed that parameters in the real quadratic family admitting an acim have positive Lebesgue measure. M. Rees proved in [31] a corresponding celebrated result for rational maps. Later, M. Benedicks and L. Carleson proved in the fundamental papers [9, 10] that the set of CE-parameters in the real quadratic family has positive Lebesgue measure. In the rational setting, a similar result was proven by the first author in [3].
Let us recall some definitions. A function $f_{c}$ is called Collet–Eckmann if there exist $C>0$ and $\gamma>0$ such that
@@
A map $f_{c}$ in the unicritical family is hyperbolic if $\omega(0)\cap{\mathcal{J}}(f_{c})=\emptyset$, where ${\mathcal{J}}(f_{c})$ is the Julia set of $f_{c}$. For $d\geq 2$, the set $\mathcal{M}_{d}$ is defined as the set of parameters $c$ such that ${\mathcal{J}}(f_{c})$ is connected (or equivalently, the set of parameters $c$ for which the critical orbit $\{f^{n}_{c}(0)\}$ is bounded). When $d=2$, the set $\mathcal{M}_{2}={\mathcal{M}}$ is the well known Mandelbrot set. (We shall also call ${\mathcal{M}}_{d}$ a Mandelbrot set.)
Our paper is concerned with (complex) perturbations of CE-maps in the unicritical family. In [32], J. Rivera-Letelier proved that Misiurewicz maps (i.e. critically non-recurrent maps) are Lebesgue density points of ${\mathbb{C}}\setminus{\mathcal{M}}_{d}$. It was generalised by the first author in [1] to rational maps. However, even if Misiurewicz maps behave quite nicely under perturbations, they are rare in the parameter space [2].
In a series of papers [17, 19, 20], J. Graczyk and G. Świa̧tek studied the geometry of the Julia set around typical maps with respect to harmonic measure on $\partial{\mathcal{M}}_{d}$. Among other things, they show that almost all maps with respect to harmonic measure are critically slowly recurrent CE-maps and density points of $\mathbb{C}$$\setminus{\mathcal{M}}$ (we refer to slow recurrence in the sense of G. Levin, F. Przytycki and W. Shen [26]).
In the other direction, A. Avila, M. Lyubich and W. Shen proved in [7] that the Lebesgue density of ${\mathcal{M}}_{d}$ at at most finitely renormalizable maps in the unicritical family, is strictly less than one (Theorem 1.4 and Remark 6.1 in [7]). Since Collet–Eckmann maps are at most finitely renormalizable, the result applies to them and consequently CE-maps cannot be density points of ${\mathcal{M}}_{d}$.
Our main theorem complements the above results as follows.
@@
One could ask if our result extends to the setting of at most finitely renormalizable maps, i.e., those considered by A. Avila, M. Lyubich and W. Shen. We remark that this is not the case, as was pointed out in [18] by J. Graczyk and G. Świa̧tek .
In the same paper just mentioned, J. Graczyk and G. Świa̧tek showed that at most finitely renormalizable maps are weak density points of ${\mathbb{C}}\setminus{\mathcal{M}}_{d}$ (again, this applies to CE-maps). Thus our theorem strengthens their result for CE-maps.
The slow recurrence condition is also imposed on the Collet–Eckmann maps in [5], where we proved a corresponding result for rational maps. We also mention an unpublished work by M. Benedicks and J. Graczyk, where they prove corresponding results for one-dimensional families (with some regularity) around typical CE-maps with respect to harmonic measure in $\partial{\mathcal{M}}_{d}$. For more general CE-maps, the dynamical constants can be drastically different. This yields quite wild dynamics, and crucial parts of the traditional methods break down.
Theorem 1.1 is also interesting from a complex analytic viewpoint. Indeed, in [33], S. Smirnov proved that for a set $E$ of angles whose complement has Hausdorff dimension 0, the external rays with angles in $E$ will land on a Collet–Eckmann parameter on the boundary of the Mandelbrot set. As a consequence, by Theorem 1.1, the Mandelbrot set is an example of a simply connected compact set where a large set of boundary points, in the above sense, has Lebesgue density zero. So from the outside, the Mandelbrot set looks like a one-dimensional object; the interior is not detectable from the outside.
Ideas of the proof. We trace the critical orbits for all parameters close to a Collet–Eckmann map and use the parameter exclusion technique developed by M. Benedicks and L. Carleson in the seminal works [9, 10]. However, since we start with arbitrary Collet–Eckmann maps, the dynamical constants, e.g. the recurrence rate, can be quite wild in general (deep central returns) and traditional methods break down. This difficulty will be overcome by introducing the idea of promotion. Roughly speaking, promotion means that almost all parameters will reach a condition arbitrarily close to the slow recurrence condition. This technique is one of the core novelties and fundamental for the result of the paper.
Acknowledgement. Partially this work was written when all the authors were participating in the program “Two Dimensional Maps” at Mittag-Leffler Institute during the spring 2023. We thank the organisers for providing a nice working place. We are also grateful to Michael Benedicks for stimulating discussions on the topic of this paper. The first author gratefully acknowledges partial funding by the Swedish Research Council. The second author gratefully acknowledges support from the Knut and Alice Wallenberg Foundation. The third author was partially supported by NSF of China (No. 12401105, 12522108), Qingdao Natural Science Foundation (No. 24-4-4-zrjj-8-jch), Shandong Provincial Natural Science Fund for Excellent Young Scientists Program (Overseas) (No. 2025HWYQ-021) and also a grant from Vergstiftelsen during his stay in Sweden.
@@ 2 Preliminaries
Throughout the paper we fix $d\geq 2$. Let $f_{c_{0}}$ be a Collet–Eckmann unicritical polynomial of degree $d$. The only critical point is 0 with the corresponding critical value being $c_{0}$. Let ${\mathcal{Q}}={\mathcal{Q}}(c_{0},\varepsilon)$ be a square of side length $\varepsilon$ in the parameter plane centred at the parameter $c_{0}$ with $\varepsilon>0$ to be sufficiently small (cf. Lemma 2.9 and Lemma 2.8). For a parameter $c\in{\mathcal{Q}}$ the corresponding map is denoted, as usual, by $f_{c}$. We study the evolution of the critical orbit for nearby parameters. Let
@@ 2.1 Recurrence, bound periods and partitions
After the bound period ends, the two critical orbits will then separate and stay away from critical points until the next return happens. A return might be deep in the sense that $\xi_{n}(c)$ is quite close to 0. To achieve our goal, we will need to disregard parameters with very deep returns. For later purpose we will define $\alpha_{n}(c)$ for a parameter $c$ as follows:
@@
$\operatorname{diam}\xi_{k}(A)\leq\begin{cases}~\,~\dfrac{\operatorname{dist}(\xi_{k}(A),0)}{(\log\operatorname{dist}(\xi_{k}(A),0))^{2}}~&\text{if}\,~\xi_{k}(A)\cap U\neq\emptyset,\\ \quad~\,~S~&\text{if}\,~\xi_{k}(A)\cap U=\emptyset.\end{cases}$ (2.6)
@@
For a partition element, we can also give a notion of returns. Let $A$ be a partition element at time k1k-1. Then $\xi_{k}(A)$ is called a return to $U$ if $\xi_{k}(A)\cap U\neq\emptyset$ and a return to $U^{\prime}$ if $\xi_{k}(A)\cap U^{\prime}\neq\emptyset$. It also may happen that $\xi_{k}(A)\cap U^{\prime}\neq\emptyset$ but $\xi_{k}(A)\cap U=\emptyset$. In this case we say that $\xi_{k}(A)$ is a pseudo-return. Now bound periods can also be defined similarly if a return occurs for a partition element.
@@ 2.2 Free periods
Let $f$ be rational satisfying the exponential shrinking property. Then there exist a neighbourhood $U^{\prime}$ of $\operatorname{Crit}(f)\cap{\mathcal{J}}(f)$, $C>0$ and $\gamma>0$ such that the following holds: Let $z\in{\mathcal{J}}(f)$ and if $f^{j}(z)\not\in U^{\prime}$ for $0\leq j\leq n-1$, then
@@
The following lemma follows from the above and is proved in [5, Lemma 3.10]. To state the result, let $\mathcal{N}_{\varepsilon_{0}}$ be a neighbourhood of the Julia set ${\mathcal{J}}(f_{c_{0}})$, where $\varepsilon_{0}$ is chosen in a way such that $U^{\prime}\subset\mathcal{N}_{\varepsilon_{0}}$. That is,
@@ 2.3 Transversality
A direct consequence of this is the following lemma proved in [4, Proposition 4.1].
@@ 3.1 Distortion, length and expansion during bound periods
Suppose that $\operatorname{dist}(0,\xi_{n}(A))\sim e^{-r}$. Then for all $z\in\xi_{n}(A)$ we have $|f_{c}(z)-c|\sim e^{-dr}$. We divide the sum into two parts $0,j-1]=[0,J]\cup[J+1,j-1]$, where $J=\lceil r/(\alpha+\overline{\gamma})\rceil$ and where $\alpha$ is the upper bound of $\alpha_{j}=\alpha_{j}(c)$ for $c\in A$, satisfying $\alpha<\overline{\gamma}-\underline{\gamma}$. Then the first sum becomes
@@ 3.2 Main Distortion Lemma
[…]
@@
For consistency we include the proof of this important lemma, although its proof is almost the same, verbatim, as the proof of Lemma 3.14 in [5] (alternatively see [4, Lemma 7.3]).
@@
where we in the sum from $q_{t}$ to n1n-1 used Lemma 2.8 (the first inequality, now with respect to $U^{\prime}$) and that $|\xi_{j}(b)|>\delta^{\prime}>\delta$ during this time.
@@ 2. $(ii)$
    $\mathcal{Q}$ is a partition element at time N1N-1, and $\xi_{N}(Q)$ either is an essential return or reaches the large scale $S$.
@@ 4. $(iv)$
    There exists a constant $K=K(\delta,\varepsilon_{1})>1$ such that for all $c\in\mathcal{Q}$ and $0\leq k\leq N-1$
@@
By the definition of a partition element we can use Lemma 3.8 repeatedly starting from the time $N_{L}$ up to N1N-1 to get the distortion claimed in (iii).
@@ 4 Promotion
By Lemma 3.9, all critical orbits for parameters in $\mathcal{Q}$} will follow the starting critical orbit up to time $N$ with bounded distortion and exponential growth of derivatives. If $\mathcal{Q}$} has grown to the large scale under $\xi_{N}$, then $\xi_{N}({\mathcal{Q}})$ covers mostly the Fatou set of $f_{0}$, and by a compactness argument, we can conclude that most parameters in $\mathcal{Q}$} will land in the Fatou set of the corresponding map under $\xi_{n}$; See the next section for more details. Hence a very large portion of parameters in $\mathcal{Q}$} are hyperbolic. Therefore, we may assume, without loss of generality, that $\xi_{N}({\mathcal{Q}})$ has not reached the large scale yet and thus is an essential return.
@@
Here, since $\alpha_{N}(c_{0})$ is bounded, $\delta_{0}=\delta_{0}(N)$ can be made arbitrarily small given $N$ large enough. We make the deletion so that the smallest collection of whole partition elements are deleted (this means that we may delete a little more than necessary). The remaining parameters are partitioned into partition elements as above. Each of these partition elements, containing $c$, is denoted by $A_{N}(c)$. Sometimes we just write $A_{N}$. However, note that $A_{N-1}={\mathcal{Q}}$ (there is no partition before time $N$).
After this time $N$ we lose this kind of dependence on the starting orbit, hence we need to control the rate of recurrence in order to not lose too much growth of the derivative.
Let $N=\nu_{0}<\nu_{1}<\cdots$ be the free returns after time $N$. Suppose that $A=A_{\nu_{j}}$ is a partition element at time $\nu_{j}$. We will inductively introduce two sequences $\{\widetilde{\alpha}_{\nu_{j}}(A)\}$ and $\{\underline{\gamma}_{\nu_{j}}(A)\}$ which capture two key features for later analysis. Firstly, parameters $c\in A$ for which $\xi_{\nu_{j}}(c)$ are at distance $e^{-\widetilde{\alpha}_{\nu_{j}}(A)\nu_{j}}$ away from the critical point will be kept (and thus has controlled approach rate to the critical point), and secondly, we do not lose a lot of the derivative growth between two free returns.
We start with the time $\nu_{0}$. By Lemma 3.9 (iv), $\xi_{k}(c)$ stays close to $\xi_{k}(c_{0})$ and thus $\alpha_{k}(c)$ is comparable to $\alpha_{k}(c_{0})$ for all $k\leq N$ and all $c\in A_{\nu_{0}}$.
Definition of $\widetilde{\alpha}_{\nu_{j}}(A)$.
@@
$\widetilde{\alpha}_{\nu_{0}}(A_{\nu_{0}})=\sup_{c\in A_{\nu_{0}}}\alpha_{\nu_{0}}(c).$
@@
Let $|\xi_{\nu_{0}}(c)|\sim e^{-r}$, for some $c\in A_{\nu_{0}}$, where $A_{\nu_{0}}\subset{\mathcal{Q}}$ is a partition element satisfying (4.1). Then $|Df_{c}^{p_{0}+1}(\xi_{\nu_{0}}(c))|\geq e^{(1-\kappa)r}$.
@@
We can make $C_{1}$ as close to $1$ as we want so that $c\in A_{\nu_{0}}$ implies $c\in{\mathcal{E}}_{n}(\underline{\gamma})$. The lemma now follows from Corollary 3.5. ∎
Note that by the definition of $\kappa$, Lemma 4.1 says that $\operatorname{diam}\left(\xi_{\nu_{0}+p_{0}}(A_{\nu_{0}})\right)\geq e^{-\kappa r}$, where $\operatorname{diam}\left(\xi_{\nu_{0}}(A_{\nu_{0}})\right)\sim e^{-r}$. Now we fix some $\kappa^{\prime}$ such that $\kappa<\kappa^{\prime}<1$. We shall now define $\widetilde{\alpha}_{\nu_{j}}(A)$, where $A$ is a partition element at the free return time $\nu_{j}$.
@@
Indeed, using the mean value theorem we find that for any $c,c^{\prime}\in A$
$\left|\frac{\alpha_{n}(c)}{\alpha_{n}(c^{\prime})}-1\right|\leq\frac{\left(\sup_{c\in A_{\nu_{j}}}e^{\alpha_{n}(c)n}\right)|\xi_{n}(c)-\xi_{n}(c^{\prime})|}{\alpha_{n}(c^{\prime})n}\sim\frac{1}{r_{j}^{3}}$
@@
$\gamma_{n}(A_{n})=\inf\limits_{c\in A_{n}}\gamma_{n}(c)\quad\text{ and }\quad\gamma_{n-1}(A_{n})=\inf\limits_{c\in A_{n}}\gamma_{n-1}(c)$
@@
$\underline{\gamma}_{\nu_{j+1}}(A)=\left(\gamma_{\nu_{j}}(A)+\gamma_{H}\frac{\ell_{j}}{\nu_{j}}\right)\frac{\nu_{j}}{\nu_{j+1}-1},$ (4.6)
@@
Note that $\underline{\gamma}_{\nu_{j}}$ assigns a number for each partition element. So, if $A_{\nu_{j}}$ and $A_{\nu_{j+1}}$ are respectively partition elements at time $\nu_{j}$ and $\nu_{j+1}$ and moreover $A_{\nu_{j+1}}\subset A_{\nu_{j}}$, we have $\underline{\gamma}_{\nu_{j}}(A_{\nu_{j+1}})=\underline{\gamma}_{\nu_{j}}(A_{\nu_{j}})$.
@@
$\geq\left(\gamma_{\nu_{j}-1}(A)\left(1-\frac{1}{\nu_{j}}\right)+(d-1)\alpha_{\nu_{j}}(c)\left(1-\frac{1}{\nu_{j}}\right)\right)\frac{\nu_{j}}{\nu_{j+1}-1}$ (4.11)
@@
$\geq\left(\gamma_{\nu_{j}-1}(A)\left(1-\frac{1}{\nu_{j}}\right)+\gamma_{H}\frac{\ell_{j}}{\nu_{j}}-\frac{C}{\nu_{j}}\right)\frac{\nu_{j}}{\nu_{j+1}-1}$ (4.13)
@@
$\frac{\gamma_{\nu_{j}-1}(A)+2C}{\nu_{j}}\leq\left((d-1)\left(1-\frac{1}{\nu_{j}}-\frac{\kappa^{\prime}}{\tilde{\kappa}}\right)+(1-\kappa)\right)\alpha_{\nu_{j}}$
for sufficiently large $\nu_{j}\geq N$, since $\alpha_{\nu_{j}}(c)\nu_{j}\sim r_{\nu_{j}}\geq\Delta$ and $\Delta$ is large. Then, using that $p_{j}\leq\nu_{j}$ by Lemma 3.2, we get $\gamma_{\nu_{j+1}}(c)\geq\underline{\gamma}_{\nu_{j+1}}(A)$. If we define $\underline{\gamma}_{k}(A)=\underline{\gamma}_{\nu_{j+1}}(A)$ for $\nu_{j}<k\leq\nu_{j+1}$, then it follows that